# On the Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ over Stable Semirings

### v1

@@ Abstract
$\text{\sf Datalog}^{\circ}$ is an extension of Datalog, where instead of a program being a collection of union of conjunctive queries over the standard Boolean semiring, a program may now be a collection of sum-sum-product queries over an arbitrary commutative partially ordered pre-semiring. $\text{\sf Datalog}^{\circ}$ is more powerful than Datalog in that its additional algebraic structure alows for supporting recursion with aggregation. At the same time, $\text{\sf Datalog}^{\circ}$ retains the syntactic and semantic simplicity of Datalog: $\text{\sf Datalog}^{\circ}$ has declarative least fixpoint semantics. The least fixpoint can be found via the naïve evaluation algorithm that repeatedly applies the immediate sequence opeator until no further change is possible.
It was shown in [13] that, when the underlying semiring is $p$-stable, then the naïve evaluation of any $\text{\sf Datalog}^{\circ}$ program over the semiring converges in a finite number of steps. However, the upper bounds on the rate of convergence were exponential in the number of ground IDB atoms.
This paper establishes polynomial upper bounds on the convergence rate of the naïve algorithm on linear $\text{\sf Datalog}^{\circ}$ programs, which is quite common in practice. In particular, the main result of this paper is that the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs under any $p$-stable semiring is $O(pn^{3})$. Furthermore, we show a matching lower bound by constructing a $p$-stable semiring and a linear $\text{\sf Datalog}^{\circ}$ program that requires $\Omega(pn^{3})$ iterations for the naïve iteration algorithm to converge. Next, we study the convergence rate in terms of the number of elements in the semiring for linear $\text{\sf Datalog}^{\circ}$ programs. When $L$ is the number of elements, the convergence rate is bounded by $O(pn\log L)$. This significantly improves the convergence rate for small $L$. We show a nearly matching lower bound as well.
@@ 1 Introduction
In order to express common recursive computations with aggregates in modern data analytics while retaining the syntactic and semantic simplicity of Datalog, [13] introduced $\text{\sf Datalog}^{\circ}$, an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative partially ordered pre-semiring (POPS). Datalog is exactly $\text{\sf Datalog}^{\circ}$ over the Boolean semiring.
Like Datalog, $\text{\sf Datalog}^{\circ}$ has a declarative least fixpoint semantics, and the least fixpoint can be found via the naïve iteration algorithm that repeatedly applies the immediate sequence opeator until no further change is possible. Moreover, its additional algebraic structure allows for common recursions with aggregations.
@@
Unlike Datalog, however, depending on the underlying semiring, the naïve evaluation of a $\text{\sf Datalog}^{\circ}$ program may not always converge in a finite number of steps. Hence, [13] studied properties of the underlying POPS for which an explicit bound on the convergence rate of the naïve algorithm can be obtained. The convergence rate of a $\text{\sf Datalog}^{\circ}$ program over a given POPS can be studied through its “core semiring”, which is where we focus our attention on in this paper. Hence, we will only study $\text{\sf Datalog}^{\circ}$ programs over semirings in this paper, referring the readers to [13] for the generality of POPS.
@@
It was showed in [13] that if the semiring $\bm{S}$ is $p$-stable111 Roughly, $\bm{S}=(S,\oplus,\otimes,0,1)$ is $p$-stable if every linear function $f:S\rightarrow S$ of the form $f(x)=a\otimes x\oplus b$ satisfies $f^{(p)}(0)=f^{(p+1)}(0)$., then the naïve evaluation of any $\text{\sf Datalog}^{\circ}$ program over $\bm{S}$ converges in a finite number of steps. The exact general upper bound on the convergence rate of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semiring is open, even for the special case of linear $\text{\sf Datalog}^{\circ}$ programs. Linear $\text{\sf Datalog}^{\circ}$ programs are quite common in practice. [^2] For example, we can express transitive closure, all-pairs-shortest-paths, or weakly connected components in linear $\text{\sf Datalog}^{\circ}$. Thus, in this paper we focus on this “easiest” case where the exact convergence rate is not known.
@@
Currently, the best known upper bound on the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings is $\sum_{i=1}^{n}(p+1)^{i}$ steps, where $n$ is the number of ground IDB atoms in the program. This bound is unsatisfactory in the following sense. The prototypical example of a $p$-stable semiring is the tropical semiring $\text{\sf Trop}_{p}^{+}$, (see Section 2 for a definition of $\text{\sf Trop}_{p}^{+}$); in this case, it is known that the naïve algorithm converges in $O(pn)$ steps for linear $\text{\sf Datalog}^{\circ}$ programs [13]. Thus, even for this “easy” case there is an exponential gap between the upper and lower bound on the convergence rate.
@@
This paper seeks to obtain tighter bounds on the convergence rate of naïve evaluation of linear $\text{\sf Datalog}^{\circ}$ programs, Bounding the convergence rates of $\text{\sf Datalog}^{\circ}$ program is of both theoretical interests and practical interests. In practice, a known upper bound on the convergence rate allows the database system to determine before hand how to verify if a $\text{\sf Datalog}^{\circ}$ program ever terminates, which is not possible with general recursion with aggregation.
@@ 1.1 Background
There is an implicit existential quantifier over the body for all variables that appear in the body but not in the head, where the domain of the existential quantifier is ADom. Thus, a Datalog program is also a collection of union of conjunctive queries (UCQs), one UCQ for each IDB. In a linear Datalog program every conjunction has at most one IDB.
@@
A Datalog program can be thought of as a function (called the immediate consequence operator, or ICO) that maps a set of ground IDB atoms to a set of ground IDB atoms. Every rule in the program is an inference rule that can be used to infer new ground IDB atoms from old. Under (ground-atom) set containment, this function has a unique least fixpoint which can be obtained via repeatedly applying the ICO until a fixpoint is reached [1]. This least fixpoint is the semantics of the given Datalog program. The algorithm is called the naïve evaluation algorithm, which converges in a polynomial number of steps in the size of the input database, given that the program is of fixed size.
Like Datalog programs, a $\text{\sf Datalog}^{\circ}$ program consists of a set of rules, where the union of conjunctive queries are now replaced with sum-sum-product queries over a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$. (As aforementioned, in this paper we focus on the core semiring of a POPS, leaving the intricacies of POPS to [13].) A Datalog program is a $\text{\sf Datalog}^{\circ}$ program over the Boolean semiring $\bm{S}=(\{\text{\sf true},\text{\sf false}\},\vee,\wedge,\text{\sf false},\text{\sf true})$.
For instance, in Example 1 the the operators $\wedge,\vee,\exists_{Z}$ are replaced by $\otimes,\oplus,\bigoplus_{Z}$. In particular, here $\oplus$ is the addition operation, and $\otimes$ is the multiplication operation, in a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$, and $\bigoplus_{Z}$ is aggregation, that is an iterative application of $\oplus$ for the domain of the variable $Z$. Each tuple in an EDB predicate or IDB predicate is associated with an element of the semiring $\bm{S}$, and the element associated with a tuple in an EDB predicate is specified in the input. Again a $\text{\sf Datalog}^{\circ}$ program is linear if every rule has no more than one IDB prediciate in its body.
@@
Every EDB and IDB atom in a $\text{\sf Datalog}^{\circ}$ program can be thought of as a function from the domain of the variables in the atom to the semiring $\bm{S}$. For instance, in (4), $T$ is a map $T:\textsf{ADom}\times\textsf{ADom}\rightarrow\mathbb{R}_{+}\cup\{\infty\}$. Since the values of the IDB “functions” are not known, it is natural to think of them as variables. For instance, in (4), $T(x,y)$ is a variable for every pair $(x,y)$. The EDB variables are known.
Definition 1.1.
Let $n$ denote the total number of IDB “variables” (i.e. the number of ground IDB atoms without the assigned value), then the immediate consequence operator (ICO) of a $\text{\sf Datalog}^{\circ}$ program is a map $f:S^{n}\rightarrow S^{n}$, defined by putting together $n$ component functions, one for each IDB variable. [^3] For instance, in (4), the ICO is $f:\mathbb{R}_{+}^{|V|^{2}}\rightarrow\mathbb{R}_{+}^{|V|^{2}}$. The EDB variables are constants in the ICO. Every component function of $f$ is a multivariate polynomial in the IDB variables. In [13], this ICO $f$ is also called the grounded $\text{\sf Datalog}^{\circ}$ program.
If the semiring is naturally ordered444 $\bm{S}$ is naturally ordered if the relation $x\preceq_{S}y$ defined as $\exists z:x\oplus z=y$ is a partial order. , then the least fixpoint of a $\text{\sf Datalog}^{\circ}$ program is the least fixed point of $f$ under the same partial order extended to $S^{n}$ componentwise. This is the least fixpoint semantics of a $\text{\sf Datalog}^{\circ}$ program. The naïve evaluation algorithm for evaluating Datalog programs extends naturally to evaluating $\text{\sf Datalog}^{\circ}$ programs: starting from $\bm{x}=0^{n}$, we repeatedly apply $f$ to $\bm{x}$ until a fixpoint is reached $\bm{x}=f(\bm{x})$. The core semiring of a POPS is naturally ordered. Thus, we can find the least fixpoint of a $\text{\sf Datalog}^{\circ}$ program by applying the naïve evaluation algorithm [13].
In this paper, we will focus on studying these ICO $f:S^{n}\to S^{n}$, and their convergence rate under the naïve iteration algorithm over a stable semiring $\bm{S}$.
@@
The following bounds on the rate of convergence of a general (multivariate) polynomial function $f:S^{n}\to S^{n}$, where $S$ is $p$-stable; this result naturally infers bounds on the convergence rate of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings.
@@
Let $f:S^{n}\to S^{n}$ be a multi-variate polynomial funciton over a $p$-stable commutative semiring $S$; then, the following holds:
  • The stability index of $f$ is at most $\sum_{i=1}^{n}(p+2)^{i}$.
  • If $f$ is linear, then its stability index is at most $\sum_{i=1}^{n}N(p+1)^{i}$.
  • If $p=0$ then the stability index of $f$ is at most $n$.
@@ 1.2 Our Results
When a $\text{\sf Datalog}^{\circ}$ program over a semiring $\bm{S}$ is linear, its ICO $f:S^{n}\rightarrow S^{n}$ is a linear function of the form $f(x)=A\otimes x\oplus b$, where $A$ is an $n$ by $n$ matrix with entries from $S$, $b$ is an $n\times 1$ column vector with entries from $S$, and the scalar multiplication and addition are from $\bm{S}$. To simplify notations, we will use $+$ and $\cdot$ to denote the operations $\oplus$ and $\otimes$ respectively, and write
$f(x)=Ax+b,$
when the semiring is clear from context.
@@
[…]
The stability index of $f$ can easily be expressed in terms of the matrix $A$ and vector $b$. Define $f^{(k)}(x)$ inductively where $f^{(1)}(x)=f(x)$ and $f^{(k)}(x)=f(f^{(k-1)}(x))$ for $k>1$. That is, $f^{(k)}$ is $k$ applications of $f$. So, defining $A^{0}=I$ where $I$ is the identity matrix, we have
@@
Thus, the linear function $f=Ax+b$ is $p$-stable if and only if $A^{(p)}b=A^{(p+1)}b$.
Definition 1.5 (Convergence Rate of a Linear $\text{\sf Datalog}^{\circ}$ Program).
  • The convergence rate for a particular linear $\text{\sf Datalog}^{\circ}$ program, with associated matrix $A$ and vector $b$, is the minimum natural number $p$ such that $A^{(p)}b=A^{(p+1)}b$
@@
Our first result is an asymptotically tight bound of $\Theta(pn^{3})$ on the rate of convergence for linear $\text{\sf Datalog}^{\circ}$ programs.
@@
In section 4 we establish a matching lower bound by finding a graph $G$ and a commutative semiring $\bm{S}$ where the calculations in the upper bound are tight.
Theorem 1.7.
There are linear $\text{\sf Datalog}^{\circ}$ programs over a $p$-stable semiring $\bm{S}$ that require $\Omega(pn^{3})$ steps to converge.
In the lower bound instance that establishes Theorem 1.7 the size of the ground set $S$ of $\bm{S}$ is of size $\Theta((p+1)^{n^{2}})$, so exponential in $n$. Thus one natural question is whether $\text{\sf Datalog}^{\circ}$ programs over semirings with subexponentially sized ground sets will converge more quickly. In section 5 we answer this question in the affirmative by showing that the rate of convergence of a $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring with $L$ elements is $O(pn\log L)$.
@@
Finally in section 6 we establish a nearly matching lower bound by finding a graph $G$ and a commutative semiring where the calculations in the upper bound are nearly tight.
Theorem 1.9.
There are linear $\text{\sf Datalog}^{\circ}$ programs over a $p$-stable commutative semiring that contains $L$ elements that require in $\Omega(pn\frac{\log L}{\log p})$ steps to converge.

### v2

@@ Abstract
$\text{\sf Datalog}^{\circ}$ is an extension of Datalog, where instead of a program being a collection of union of conjunctive queries over the standard Boolean semiring, a program may now be a collection of sum-product queries over an arbitrary commutative partially ordered pre-semiring. $\text{\sf Datalog}^{\circ}$ is more powerful than Datalog in that its additional algebraic structure alows for supporting recursion with aggregation. At the same time, $\text{\sf Datalog}^{\circ}$ retains the syntactic and semantic simplicity of Datalog: $\text{\sf Datalog}^{\circ}$ has declarative least fixpoint semantics. The least fixpoint can be found via the naïve evaluation algorithm that repeatedly applies the immediate consequence operator until no further change is possible.
It was shown in [12] that, when the underlying semiring is $p$-stable, then the naïve evaluation of any $\text{\sf Datalog}^{\circ}$ program over the semiring converges in a finite number of steps. However, the upper bounds on the rate of convergence were exponential in the number $n$ of ground IDB atoms.
This paper establishes polynomial upper bounds on the convergence rate of the naïve algorithm on linear $\text{\sf Datalog}^{\circ}$ programs, which is quite common in practice. In particular, the main result of this paper is that the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs under any $p$-stable semiring is $O(pn^{3})$. Next, we study the convergence rate in terms of the number of elements in the semiring for linear $\text{\sf Datalog}^{\circ}$ programs. When $L$ is the number of elements, we show that the convergence rate is bounded by $O(pn\log L)$. This significantly improves the convergence rate for small $L$.
keywords:
Datalog, convergence rate, semiring
@@ 1 Introduction
In order to express common recursive computations with aggregates in modern data analytics while retaining the syntactic and semantic simplicity of Datalog, [12] introduced $\text{\sf Datalog}^{\circ}$, an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative partially ordered pre-semiring (POPS). Datalog is exactly $\text{\sf Datalog}^{\circ}$ over the Boolean semiring. Like Datalog, $\text{\sf Datalog}^{\circ}$ has a declarative least fixpoint semantics, and the least fixpoint can be found via the naïve iteration algorithm that repeatedly applies the immediate consequence operator until no further change is possible. Moreover, its additional algebraic structure allows for common recursions with aggregations.
@@
However unlike Datalog, the naïve evaluation of a $\text{\sf Datalog}^{\circ}$ program may not always converge in a finite number of steps. The convergence of a $\text{\sf Datalog}^{\circ}$ program over a given POPS can be studied through its “core semiring”, which is where we focus our attention on in this paper. This paper will only study $\text{\sf Datalog}^{\circ}$ programs over commutative semirings, referring the readers to [12] for the generality of POPS.
It is known that the commutative semirings for which the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is guaranteed to converge are exactly those semirings that are stable [12]. A semiring is $p$-stable if the number of iterations required for any one-variable recursive linear $\text{\sf Datalog}^{\circ}$ program to reach a fixed point is at most $p$, and a semiring is stable if there exists a $p$ for which it is $p$-stable. Further, every non-stable semiring has a simple (linear) $\text{\sf Datalog}^{\circ}$ program with one variable with the property that the iterative evaluation of this program over that semiring will not converge. Thus it is natural to concentrate on $\text{\sf Datalog}^{\circ}$ programs over stable semirings. Previously, the best known upper bound on the convergence rate, which is the number of iterations until convergence, is $\sum_{i=1}^{n}(p+2)^{i}=\Theta(p^{n})$ steps, where $n$ is the number of ground atoms for the IDB’s that ever have a nonzero value at some point in the iterative evaluation of the $\text{\sf Datalog}^{\circ}$ program, and the underlying semiring is $p$-stable. In contrast there are no known lower bounds that show that iterative evaluation requires an exponential (in the parameter $n$) number of steps to reach convergence.
@@
The exact general upper bound on the convergence rate of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings is open, even for the special case of linear $\text{\sf Datalog}^{\circ}$ programs. Linear $\text{\sf Datalog}^{\circ}$ programs are quite common in practice. [^1] For example, we can express transitive closure, all-pairs-shortest-paths, or weakly connected components in linear $\text{\sf Datalog}^{\circ}$. Thus, in this paper we focus on this “easiest” case where the exact convergence rate is not known.
@@
Currently, the best known upper bound on the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings is $\sum_{i=1}^{n}(p+1)^{i}$ steps. This bound is unsatisfactory in the following sense. The prototypical example of a $p$-stable semiring is the tropical semiring $\text{\sf Trop}_{p}^{+}$, (see Section 2 for a definition of $\text{\sf Trop}_{p}^{+}$); in this case, it is known that the naïve algorithm converges in $O(pn)$ steps for linear $\text{\sf Datalog}^{\circ}$ programs [12]. These results leave open the possibility that the convergence rate of the naïve algorithm on linear $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings could be exponentially smaller than the best known guarantee.
@@
This paper seeks to obtain tighter bounds on the convergence rate of naïve evaluation of linear $\text{\sf Datalog}^{\circ}$ programs. As the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is a reasonably natural and important algorithm/process, bounding the running time of this process is of both theoretical interests and practical interests. In practice, a known upper bound on the convergence rate allows the database system to determine before hand an upper bound on the number of iterations that will be required to evaluation a particular $\text{\sf Datalog}^{\circ}$ program.
@@ 1.1 Background
There is an implicit existential quantifier over the body for all variables that appear in the body but not in the head, where the domain of the existential quantifier is ADom. In a linear Datalog program every conjunction has at most one IDB.
@@
A Datalog program can be thought of as a function (called the immediate consequence operator, or ICO) that maps a set of ground IDB atoms to a set of ground IDB atoms. Every rule in the program is an inference rule that can be used to infer new ground IDB atoms from old ones. For a particular EDB instance, this function has a unique least fixpoint which can be obtained via repeatedly applying the ICO until a fixpoint is reached [1]. This least fixpoint is the semantics of the given Datalog program. The algorithm is called the naïve evaluation algorithm, which converges in a polynomial number of steps in the size of the input database, given that the program is of fixed size.
Like Datalog programs, a $\text{\sf Datalog}^{\circ}$ program consists of a set of rules, where the unions of conjunctive queries are now replaced with sum-product queries over a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$; see Section 2. Each rule has the form:
[…]
where sum ranges over the active ADom of the variables not in $X_{0}$. Further each ground atom in an EDB predicate or IDB predicate is associated with an element of the semiring $\bm{S}$, and the element associated with a tuple in an EDB predicate is specified in the input. The EDBs form the input database, and the ground atoms in the IDB’s that have an associated semiring value that is nonzero represent the output instance computed by the $\text{\sf Datalog}^{\circ}$ program. Note that in a Datalog program the ground atom present in the input or output databases can be thought of as those that are associated with the element $1$ in the standard Boolean semiring. A Datalog program is a $\text{\sf Datalog}^{\circ}$ program over the Boolean semiring $\bm{S}=(\{\text{\sf true},\text{\sf false}\},\vee,\wedge,\text{\sf false},\text{\sf true})$. Again a $\text{\sf Datalog}^{\circ}$ program is linear if every rule has no more than one IDB prediciate in its body.
@@
A $\text{\sf Datalog}^{\circ}$ program can be thought of as an immediate consequence operator (ICO). To understand how the ICO works in $\text{\sf Datalog}^{\circ}$, consider a rule with head $R$ and let $A$ be a ground atom for $R$ with associated semiring value $y$, and assume that for $A$ the body of this rule evaluates to the semiring element $x$. As a result of this, the ICO associates $A$ with with $x\oplus y$. Note that the functioning of the $\text{\sf Datalog}^{\circ}$ ICO, when the semiring is the standard Boolean semiring, is identical to how the Datalog ICO functions. The iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program works in rounds/steps, where initially the semiring element 0 is associated with each possible ground atom in an IDB, and on each round the ICO is applied to the current state. So in the context of the $\text{\sf Datalog}^{\circ}$ program in Example 1, if $(a,b)$ was a ground atom in $T$ with associated semiring element $x$, meaning that the shortest known directed path from vertex $a$ to vertex $b$ has length $x$, and $(b,c)$ was a ground atom in $E$ with associated semiring element $y$, meaning that there is a directed edge from $b$ to $c$ with length $y$, then the ICO would make the semiring element associated with $A$ the minimum (as minimum is the addition operation in the tropical semiring) of its current value and $x+y$ (as normal additional is the multiplication operation in the tropical semiring).
Since the final associated semiring values of the ground atoms in an IDB are not initially known, it is natural to think of them as (IDB) variables. Then the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$program is a map $\bm{f}:S^{n}\rightarrow S^{n}$, where $n$ is the number of ground atoms for the IDB’s that ever have a nonzero value at some point in the iterative evaluation of the $\text{\sf Datalog}^{\circ}$ program. For instance, in (5), there would be one variable for each pair $(x,y)$ of vertices where there is a directed path from $x$ to $y$ in the graph. So the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$ program has the following form:
$-\displaystyle X_{1}\text{ :- }$- $f_{1}(X_{1},\ldots,X_{n})$
$\ldots$ (6)
$-\displaystyle X_{n}\text{ :- }$- $f_{n}(X_{1},\ldots,X_{n})$
where the $X_{i}$’s are the IDB variables, and $f_{i}$ is the component of $\bm{f}$ corresponding to the IDB variable $X_{i}$. Note that each component function $f_{i}$ is a multivariate polynomial in the IBD variables of degree at most the maximum number of terms in any product in the body of some rule in the $\text{\sf Datalog}^{\circ}$ program. After $q$ iterations of the iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program, the semiring value associated with the ground atom corresponding to $X_{i}$ will be:
${f}_{i}^{(q)}(\bm{0})$ (7)
@@
The following bounds on the rate of convergence of a general (multivariate) polynomial function $f:S^{n}\to S^{n}$, where $\bm{S}$ is $p$-stable; this result naturally infers bounds on the convergence rate of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings.
@@
The convergence rate of a $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring is at most $\sum_{i=1}^{n}(p+2)^{i}$. Further, the convergence rate is at most $\sum_{i=1}^{n}(p+1)^{i}$ if the $\text{\sf Datalog}^{\circ}$ program is linear. Finally, the convergence rate of a $\text{\sf Datalog}^{\circ}$ program is at most $n$ if $p=0$.
@@ 1.2 Our Results
When a $\text{\sf Datalog}^{\circ}$ program over a semiring $\bm{S}$ is linear, its ICO $f:S^{n}\rightarrow S^{n}$ is a linear function of the form $f(x)=A\otimes x\oplus b$, where $A$ is an $n$ by $n$ matrix with entries from $S$, $b$ is an $n\times 1$ column vector with entries from $S$, and the scalar multiplication and addition are from $\bm{S}$. To simplify notations, we will use $+$ and $\cdot$ to denote the operations $\oplus$ and $\otimes$ respectively. Further, following the convention, we may omit $\otimes$ if it is clear from the context. Then, we have
$f(x)=Ax+b$
@@
$A_{(u,v),(u,w)}=\begin{cases}E(v,w)&\text{if }u\neq v\\ 0&\text{otherwise}\end{cases}$ (8)
The stability index of $f$ can easily be expressed in terms of the matrix $A$ and vector $b$. Letting $A^{0}=I$ where $I$ is the identity matrix, we have
@@
Thus, the linear function $f=Ax+b$ is $p$-stable if and only if $A^{(p)}b=A^{(p+1)}b$. Using the simple form of $f$ for the linear case, we can rewrite Definition 1.2 into the following simpler form. {observation}[Convergence Rate of a Linear $\text{\sf Datalog}^{\circ}$ Program]
  • The convergence rate for a particular linear $\text{\sf Datalog}^{\circ}$ program, with associated matrix $A$ and vector $b$, is the minimum natural number $p$ such that $A^{(p)}b=A^{(p+1)}b$.
@@
Our first result is a bound of $O(pn^{3})$ on the rate of convergence for linear $\text{\sf Datalog}^{\circ}$ programs.
@@
In the conference version of this paper [10], and the original arxiv version of the paper, we claimed a matching lower bound. That is we claimed that for any $p,n\geq 1$, there is a linear $\text{\sf Datalog}^{\circ}$ program over a $p$-stable semiring $\bm{S}$ that requires $\Omega(pn^{3})$ steps to converge. It was subsequently pointed out that our claim was incorrect [6]. Still in Section 4 we briefly discuss our lower bound attempt, and the underlying conceptual error that led to our mistake, as we believe that it is instructive.
A natural question is whether $\text{\sf Datalog}^{\circ}$ programs over semirings with small sets will converge more quickly. In section 5 we answer this question in the affirmative by showing that the rate of convergence of a $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring with $L$ elements is $O(pn\log L)$.
@@
In the conference version of this paper [10], and the original arxiv version of the paper, we claimed a nearly matching lower bound. We now believe that this lower bound is also not correct, for reasons roughly similar to those that led to our incorrect $\Omega(pn^{3})$ lower bound.